OpenStudy (anonymous):

Tuan simplified the following expression: (b7)(b-3)/b2 (The 7,-3, and 2 are exponents.) His final answer was in the form bn. If he simplified the expression correctly, what is the value of n, the exponent on b?

OpenStudy (campbell_st):

well you need to know the index law for multiplication. same base add the powers \[x^a \times x^b = x^{a + b}\] so start by simplifying \[b^7 \times b^{-3} \] just add the powers. next the index law for division is... same base, subtract the powers. \[\frac{x^c}{x^d} = x^{c - d} \] you will need this law for the 2nd part \[\frac{b^{7 - 3}}{b^2} \] hope this helps

OpenStudy (anonymous):

campbell_st do u know the answer? i have tried and tried

OpenStudy (campbell_st):

so this will give the answer... \[b^{7 + (-3) - 2} \] just add the powers.